Using retro synthesis what alkyl halide and ketone were used to produce 3-ethyl-3-hexanol. The two starting materials cannot be 1-bromopropane and 3-pentanone.

Calculate the percent yield of the final product which weighed 0.16g. The starting weight of Magnesium was 0.612g and starting weight of Ammonium Chloride was 0.6012g.
Answer #1

2)

    Mg + 2 NH4Cl ------------ MgCl2 + 2 NH3 + H2

1 mole    2 mole                  1 mole

mass of Mg= 0.612 g

molar mass of Mg = 24.0gram/mole

number of moles of Mg = mass/molar mass = 0.612/24.0= 0.0255 mole

mass of NH4Cl = 0.6012 grams

molar mass of NH4Cl = 53.49 gram/mole

number of moles of NH4Cl = 0.6012/53.49 = 0.0112 moles

according to equation

1 mole of Mg = 2 mole of NH4Cl

0.0255 moles of Mg = ?

                                = 0.0255x2/1= 0.051 mole of Nh4Cl

we need 0.051 moles of NH4Cl. but we have 0.0112 mole of NH4Cl. so , NH4Cl is completed first in the reaction .

Hence NH4Cl is limiting reagent.

according to euation

2 mole of NH4Cl = 1 mole ofMgCl2

0.0112 mole of NH4Cl = ?

                                        = 0.0112x1/2= 0.0056 mole

number of moles of MgCl2 formed = 0.0056 moles

molar mass of MgCl2 = 95.2 gram/mole

mass of 0.0056 moles ofMgCl2 = 0.0056 x 95.2 = 0.53312 grans

Theoritical yield of MgCl2 = 0.533 grams

Actual yield = 0.16 gram

Percent yield = actual yield/theoritical yield x100 = 0.16/0.533 x100=30.018 %

Percent yield = 30.02%.